OpenStudy (anonymous):

let f(x) = - 10x + 4. find F(1/2) =

OpenStudy (ivettef365):

substitute 1/2 for x then you have -10(1/2)+4

OpenStudy (ivettef365):

Therefore it is -10/4 + 4 -5/2+4 = 3/2

OpenStudy (ivettef365):

sorry wrong -10/2 + 4 = -5 + 4 = =1

OpenStudy (ivettef365):

answer -1